How do you solve the quadratic equation
16x^2-21=0

Answer by stanbon(75887)

How do you solve the quadratic equation
16x^2-21=0
----
Factor:
(4x-sqrt(21))(4x+sqrt(21)) = 0
-----
x = sqrt(21)/4 or x = -sqrt(21)/4
================
Cheers,
Stan H.
================